# xTrace: A Facial Expressive Behaviour Analysis Tool for Continuous Affect Recognition

### v1

@@ Abstract.
Recognising expressive behaviours in face videos is a long-standing challenge in Affective Computing. Despite significant advancements in recent years, it still remains a challenge to build a robust and reliable system for naturalistic and in-the-wild facial expressive behaviour analysis in real time. This paper addresses two key challenges in building such a system: (1). The paucity of large-scale labelled facial affect video datasets with extensive coverage of the 2D emotion space, and (2). The difficulty of extracting facial video features that are discriminative, interpretable, robust, and computationally efficient. Toward addressing these challenges, we introduce xTrace, a robust tool for facial expressive behaviour analysis and predicting continuous values of dimensional emotions, namely valence and arousal, from in-the-wild face videos.
@@
To address challenge (1), our affect recognition model is trained on the largest facial affect video data set, containing  450k videos that cover most emotion zones in the dimensional emotion space, making xTrace highly versatile in analysing a wide spectrum of naturalistic expressive behaviours. To address challenge (2), xTrace uses facial affect descriptors that are not only explainable, but can also achieve a high degree of accuracy and robustness with low computational complexity. The key components of xTrace are benchmarked against three existing tools: MediaPipe, OpenFace, and Augsburg Affect Toolbox. On an in-the-wild validation set composed of 50k videos, xTrace achieves 0.86 mean CCC and 0.13 mean absolute error values. We present a detailed error analysis of affect predictions from xTrace, illustrating (a). its ability to recognise emotions with high accuracy across most bins in the 2D emotion space, (b). its robustness to non-frontal head pose angles, and (c). a strong correlation between its uncertainty estimates and its accuracy.
Facial Expression Analysis, Dimensional Emotion Recognition, Uncertainty-Aware Affect Estimation
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXisbn:
@@
Figure 1. Extensive predictive range of xTrace for continuous affect recognition within the 2D Emotion space
1. Introduction
@@
Automatic recognition of facial expressive behaviour is a critical component in enabling empathetic Human-Computer Interactions. During the last three decades, several efforts have been made to build tools for recognising apparent affect or emotion from face videos (Pantic and Rothkrantz, 2000; Valstar and Pantic, 2006; Sanchez et al., 2021; Kollias et al., 2025). Despite the significant progress made so far, robust facial affect recognition in naturalistic conditions is still not a solved problem. Why is it challenging to build such a robust tool capable of handling real-world complexities?
@@
Affective states and their nuances encountered in everyday naturalistic interactions go beyond the seven basic emotions (Ekman, 1999): neutral, happy, angry, sad, disgust, fear, and surprise. Most of the existing methods in the literature (Tian et al., 2011; Jiang et al., 2020; Wang et al., 2022) and the widely used tools for facial expression analysis (e.g. DeepFace [^1] https://github.com/serengil/deepface) largely focus on the seven basic emotions, severely limiting their applicability to real-world scenarios. Most importantly, due to its discrete nature, the basic emotion model fails to capture the richness of emotional expressions encountered in naturalistic interactions. The dimensional model of continuous emotions (Russell, 1980) is considered one solution to address this limitation. As shown in Figure 1, the dimensional model has the ability to describe a wide range of emotions. It does so by representing the affective states as points living in a two-dimensional Euclidean space constructed by the dimensions: Valence (positive vs. negative) and Arousal (calm vs. excited). Thus, the dimensional emotion model makes it possible to estimate facial affect as continuous values in the 2D emotion space and time.
@@
The dimensional model has received a lot of interest in Affective Computing in recent years (Nicolaou et al., 2011; Gunes et al., 2011; Toisoul et al., 2021), however, its application in real-world systems is still very limited. One main reason for this is the paucity of large-scale in-the-wild affect-labelled video datasets for facial expressive behaviour analysis. Existing video datasets are relatively small in scale, with limited coverage of affective states in the 2D emotion space. For example, the Aff-wild-2 dataset (Kollias and Zafeiriou, 2018) has almost no data for the low-arousal quadrants (see Figure 3). As a result, the models trained and validated on such datasets, despite showing impressive accuracy levels, offer limited practical use in recognising a wide range of apparent emotions. In this work, we introduce a novel facial expressive behaviour analysis tool, xTrace, a versatile tool for recognising expressed facial affect in real time. Towards addressing the aforementioned limitations of the existing video affect datasets, we use the largest in-the-wild facial video affect dataset composed of  450k in-the-wild face videos to train the continuous affect recognition model in xTrace. As a result, this model shows wide predictive range of affective states in the dimensional emotion space, as shown in Figure 1.
@@
  2. (2)
    The ability to leverage long-range temporal video context for reliable emotion predictions
  3. (3)
    The use of explainable and interpretable features for traceability and transparency to mitigate model bias risks
  4. (4)
    Robustness to partial facial occlusions (induced by non-frontal head pose angles and external objects such as glasses, facial hair, etc.)
  5. (5)
    The ability to output predictive uncertainty estimates for the benefit of downstream tasks
  6. (6)
    Invariance to input image sensors (RGB vs. Near-Infra Red)
The affect recognition model in xTrace is extensively evaluated w.r.t. its input feature accuracy, its output accuracy variations across the emotion space with varying levels of granularity, its robustness to occlusions, and the quality of its predictive uncertainty estimates. The key components of xTrace, 2D face alignment, Action Unit (AU) recognition, and affect recognition, are extensively benchmarked against existing tools such as MediaPipe (Lugaresi et al., 2019), OpenFace (Tadas et al., 2018) and Augsburg Affect Toolbox (Mertes et al., 2024). The evaluation results show that xTrace outperforms these tools by significant margins on in-the-wild validation datasets.
@@
  • Shows minimal accuracy degradation in the presence of facial occlusions induced by non-frontal head pose angles and external objects.
  • Uses explainable and interpretable facial affect descriptors: uncertainty-aware facial landmarks and AU intensities.
@@
2. System Architecture
Figure 2. The end-to-end pipeline of xTrace for temporal valence and arousal estimation from face videos
@@ 2. System Architecture
2.1. Key System Design Choices
@@
Affect Feature Descriptors Broadly, there are two categories of facial affect feature sets: High-Level Descriptors (HLDs) and Deep Visual Descriptors. The former set of descriptors is derived from pre-defined facial geometry information and some standard facial actions. The latter set of features are inferred directly from cropped face images and/or image sequences using an end-to-end deep neural network trained using large amounts of affect-labelled data. A commonly used set of HLDs in the literature on facial affect analysis are facial landmarks (Jin and Tan, 2017) and facial action units (Ekman, 1977). With the advent and success of deep representation learning in the past ten years, various deep visual features have been explored for facial affect recognition tasks. EmoNet (Toisoul et al., 2021) is a state-of-the-art approach based on deep visual features jointly learnt deep visual features for landmark regression and affect recognition.
@@
As facial affect descriptors in xTrace, we chose to use HLDs, particularly 2D facial landmark features, their confidence scores and AU intensities over deep visual features for the following reasons:
@@
Explainability and Interpretability. Unlike deep visual features, HLDs are easy to explain and interpret given that they are standardised affect descriptors strongly grounded in localised facial appearance changes. For example, AU intensity features capture atomic movements of facial muscles and their temporal dynamics. When used as input features, AU features render transparency which deep visual features lack due to their black box and hierarhcically compostional nature. For example, humans easily interpret the meaning of specific facial actions like a brow-raise (AU1, AU2) or a smile (AU12), whereas deep visual features often lack such direct human interpretability. Given this advatnage with HLDs, we can trace the affect recognition model’s emotion predictions back to its corresponding input features, and precisely attribute them to visible appearance changes in different face regions. This traceability is particularly helpful in characterising the model’s failure cases and improving the model’s accuracy and robustness.
@@
Efficient use of labelled affect data. When training an end-to-end model for affect recognition from face videos, the labelled data are used for supervision for two subtasks: learning per-frame visual features from face image pixels and learning temporal dynamics that are relevant to affect prediction. The former task involves distilling facial expression visual features that are robust to nuisance factors such as changes in head pose, lighting, occlusions, etc. To achieve such a high degree of robustness, it requires large amounts of affect-labelled data or sophisticated pre-training techniques (e.g., SwaV (Caron et al., 2020)) or joint-learning methods (e.g., EmoNet (Toisoul et al., 2021)). To achieve high affect recognition accuracy, it also requires large amounts of temporal variations in affect labels in training sets, to learn temporal dynamics of the facial features. Achieving both these objectives requires enormously large amounts of affect-labelled data, obtaining which is prohibitively expensive due to the requirement of trained human annotators for this task. To efficiently use the labelled data, in this work we chose HLDs whose models are trained to be robust to changes in head-pose angles, lighting, and occlusions. Thus, in the proposed system all the affect-labelled data gets mainly used for learning the temporal dynamics of affect labels.
Leveraging long-range temporal context. Considering how crucial temporal elements are in facial emotions, the use of temporal context is known to be critical to improving the accuracy of affect recognition systems (Meng et al., 2015; Tellamekala and Valstar, 2019; Sanchez et al., 2021). Due to their high computational complexity and large feature dimensionality, end-to-end architectures used for deep visual feature extraction typically use short temporal sequences at training time, thus limiting the amount of temporal context they can leverage effectively. In contrast, given the HLDs’ low-dimensionality and the computationally light-weight regression model used, the affect recognition model in xTrace can make use of long-range temporal context (e.g. 4 to 5 seconds) for predicting VA values with better accuracy levels.
Low computational complexity and inference latency. The HLD-based affect recognition model is significantly less complex in its computational requirements, as it focuses mainly on temporal dynamics and offloads feature extraction to the ‘backbone’ models, which are landmark regression and AU estimation in xTrace. When these backbone models are heavily optimised for low inference complexity and latency, the overall system is much lighter compared to an end-to-end model such as EmoNet. It is also worth noting that unlike the end-to-end architectures, the architecture of xTrace makes it much easier to distribute its modules to different compute processing units on low-end embedded devices.
@@
Agnostic to low-level image characteristics. This is a critical requirement to make an affect recognition system invariant to changes in image acquisition process. For example, an end-to-end model trained on RGB images is likely to fail when applied to Near-Infrared (NIR) images of arbitrary wavelengths. This could be due to the dependence of deep visual features on low-level pixel statistics. Due to this limitation, an additional model fine-tuning effort may be required when there is a change in the image acquisition steps. In xTrace, by making the models responsible for HLDs robust to such low-level pixel patterns, the affect recognition model is made agnostic to any low-level image characteristics.
2.2. Affect Regression Outputs
@@
Time- and value-continuous predictions of affect. The affect regression model in xTrace is a sequence-to-sequence mapping model that takes a sequence of HLDs per frame as input and outputs a sequence of valence and arousal estimates. By mapping each frame to its corresponding continuous-valued vector in the 2D emotion space, this fine-grained approach allows us to make full use of the emotion representation capacity of the dimensional affect model. Also, by outputting per-frame affect vectors, this model captures transient and sudden changes in affective states over time. This level of spatio-temporal granularity broadens the applicability of xTrace for a broad spectrum of behavioural analysis tasks.
@@
Sampling-free predictive uncertainty estimates of affect. In addition to the point-wise affect predictions, xTrace also outputs three types of predictive uncertainty estimates of affect: Epistemic, Aleatoric, and cumulative (Hüllermeier and Waegeman, 2021). Most importantly, unlike other existing methods (Kendall and Gal, 2017; Tellamekala et al., 2022), it does so in a sampling-free manner, i.e., without requiring multiple forward passes through the model at inference time. This ability is important in obtaining uncertainty estimates with limited computational resources at inference time. Predictive uncertainties are essential for downstream applications to decide when they can or cannot trust the affect predictions. The three types of affect uncertainty output from xTrace have specific use cases. Epistemic uncertainty or model uncertainty allows the system to flag and collect out-of-distribution inputs and guide training data collection in later model update iterations. This feature is critical to continuously improve the coverage, accuracy, and robustness of the system over time. The aleatoric or data uncertainty captures the stochastic noise inherent in the input video. For example, when there is strong illumination on the face making it difficult to read the expression, the aleatoric uncertainty is expected to be high. When implemented in real-world applications, having access to this component of uncertainty can guide any potential automated interventions to improve the quality of face video. The cumulative uncertainty captures both epistemic and aleatoric components and it can be made available to a downstream task to decide how reliable the affect predictions are for a given input.
2.3. Module Descriptions
@@
Uncertainty-Aware Landmark Detection and Tracking. This module takes a video frame as input and outputs (a). a binary label for face shape validity (b). a 2D array of 68 landmark coordinates in iBUG format (Sagonas et al., 2016) and (c). a 1D array of landmark-wise uncertainty estimates. This module uses an edge-friendly CNN architecture with three output heads to jointly predict these three outputs. This CNN model is trained to be robust to in-the-wild facial appearances, on an internal curated face image dataset containing  650k RGB and NIR images labelled with 68 landmarks and their visibility scores.
Note that the face-shape validity output is learned in an unsupervised fashion, whereas the landmark-wise uncertainty values are learned using explicit supervision. If the predicted face shape validity label is false for the current frame in the buffer, its corresponding affect features (both landmarks and AUs) are set to zeros and the next video frame in the buffer gets processed. If not, the predicted landmarks are used to preprocess the input to the AU intensity estimation module.
This module first extracts the face crop from the video frame using the bounding box computed from the previous frame’s landmark coordinates. In case there is no previous frame or its predicted face shape validity is False, the bounding box is obtained from a face detection model heavily optimised for low inference latency and high detection accuracy on the edge devices.
@@
AU Intensity Estimation model takes an image of similarity-transformed face crop as input and outputs a 15-dimensional AU intensity vector. The predicted AU intensity outputs are continuous values in the range [0,5]; 0 indicates no activation and 5 indicates maximum activation. Similar to the landmark detection and tracking model, this module also uses an edge-friendly CNN architecture for compute-efficient model inference. By training on an internally curated dataset that contains  2.4 million in-the-wild facial appearances, the AU model is made robust to the presence of partial occlusions and non-uniform illumination.
@@
3. Video Datasets for Facial Affect Recognition

### v2

@@ Abstract
Recognising expressive behaviours in face videos is a long-standing challenge in Affective Computing. Despite significant advancements in recent years, it still remains a challenge to build a robust and reliable system for naturalistic and in-the-wild facial expressive behaviour analysis in real time. This paper addresses two key challenges in building such a system: (1). The paucity of large-scale labelled facial affect video datasets with extensive coverage of the 2D emotion space, and (2). The difficulty of extracting facial video features that are discriminative, interpretable, robust, and computationally efficient. Toward addressing these challenges, this work introduces xTrace, a robust tool for facial expressive behaviour analysis and predicting continuous values of dimensional emotions, namely valence and arousal, from in-the-wild face videos.
@@
To address challenge (1), the proposed affect recognition model is trained on the largest facial affect video data set, containing $\sim$450k videos that cover most emotion zones in the dimensional emotion space, making xTrace highly versatile in analysing a wide spectrum of naturalistic expressive behaviours. To address challenge (2), xTrace uses facial affect descriptors that are not only explainable, but can also achieve a high degree of accuracy and robustness with low computational complexity. The key components of xTrace are benchmarked against three existing tools: MediaPipe, OpenFace, and Augsburg Affect Toolbox. On an in-the-wild benchmarking set composed of $\sim$50k videos, xTrace achieves 0.86 mean Concordance Correlation Coefficient (CCC) and on the SEWA test set it achieves 0.75 mean CCC, outperforming existing SOTA by $\sim$7.1%. We present a detailed error analysis of affect predictions from xTrace, illustrating (a). its ability to predict emotions with high accuracy across most bins in the 2D emotion space, (b). its robustness to non-frontal head pose angles, and (c). a strong correlation between its uncertainty estimates and its accuracy.
@@
Figure 1: Extensive predictive range of xTrace for continuous affect recognition within the 2D Emotion space
I Introduction
@@
Automatic recognition of facial expressive behaviour is a critical component in enabling empathetic Human-Computer Interactions. During the last three decades, several efforts have been made to build tools for recognising apparent affect or emotion from face videos [20, 31, 23, 11]. Despite the significant progress made so far, robust facial affect recognition in naturalistic conditions is still not a solved problem. Why is it challenging to build such a robust tool capable of handling real-world complexities?
@@
Affective states and their nuances encountered in everyday naturalistic interactions go beyond the seven basic emotions [3]: neutral, happy, angry, sad, disgust, fear, and surprise. Most of the existing methods in the literature [29, 6, 32] and the widely used tools for facial expression analysis (e.g. DeepFace [^2] https://github.com/serengil/deepface) largely focus on the seven basic emotions, severely limiting their applicability to real-world scenarios. Most importantly, due to its discrete nature, the basic emotion model fails to capture the richness of emotional expressions encountered in naturalistic interactions. The dimensional model of continuous emotions [21] is considered one solution to address this limitation. As shown in Figure 1, the dimensional model has the ability to describe a wide range of emotions. It does so by representing affective states as points living in a two-dimensional Euclidean space constructed by the dimensions: Valence (positive vs. negative) and Arousal (calm vs. excited). Thus, the dimensional emotion model makes it possible to represent facial affect as continuous values in the 2D emotion space and time.
@@
The dimensional model has received a lot of interest in Affective Computing in recent years [19, 4, 30], however, its application in real-world systems is still very limited. One main reason for this is the paucity of large-scale in-the-wild affect-labelled video datasets for facial expressive behaviour analysis. Existing video datasets are relatively small in scale, with limited coverage of affective states in the 2D emotion space. For example, the Aff-wild-2 dataset [12] has almost no data for the low-arousal quadrants (see Figure 3). As a result, the models trained and validated on such datasets, despite showing impressive accuracy levels, offer limited practical use in recognising a wide range of apparent emotions. In this work, we introduce a novel facial expressive behaviour analysis tool, xTrace, a versatile tool for recognising expressed facial affect in real time. Towards addressing the aforementioned limitations of the existing video affect datasets, we use the largest in-the-wild facial affect dataset composed of $\sim$450k in-the-wild videos to train the continuous affect recognition model in xTrace. Thus, this model shows wide predictive range of affective states in the dimensional emotion space, as shown in Figure 1.
@@
  2. 2.
    The use of interpretable affect features for explainability
  3. 3.
    Robustness to partial facial occlusions induced by non-frontal head pose angles in unconstrained settings
  4. 4.
    The ability to output predictive uncertainty estimates
  5. 5.
    Invariance to image modality (RGB vs. Near-Infra Red)
The affect recognition model in xTrace is extensively evaluated w.r.t. its input feature accuracy, its output accuracy variations across the emotion space with varying levels of granularity, its robustness to occlusions, and the quality of its predictive uncertainty estimates. The key components of xTrace, 2D face alignment, Action Unit (AU) recognition, and affect recognition, are extensively benchmarked against existing tools such as MediaPipe [16], OpenFace [25] and Augsburg Affect Toolbox [17]. The evaluation results show that xTrace outperforms these tools by significant margins on in-the-wild benchmarking datasets.
xTrace is heavily optimised for on-device processing in real-time (>=>=30FPS) using only $\sim$433.5M FLOPS per frame, with cross-platform compatibility (x86 and ARM devices). Its executable runs on Windows, Mac and Linux, and it can be made available for research purposes through an academic license request process for responsible use.
@@
  • Achieves SOTA performance on the SEWA test set by outperforming the best existing method by $\sim$7.1%.
  • Uses explainable and interpretable facial affect descriptors: uncertainty-aware landmarks and AU intensities.
  • Shows minimal accuracy degradation in the presence of facial occlusions induced by non-frontal head pose.
@@
II System Architecture
Figure 2: The end-to-end pipeline of xTrace for temporal valence and arousal estimation from face videos
@@ II System Architecture
II-A Key System Design Choices
@@
Facial Affect Descriptors. Models trained to recognise affect from face videos use broadly two types of features extracted from images: Low-Level Descriptors (LLDs) and deep visual features. The former type of features are derived from spatio-temporal changes in 2D or 3D facial geometry and facial muscle actions. The latter set of features are the feature embeddings directly extracted from cropped face images and/or image sequences using a deep neural network trained using large amounts of affect-labelled data. The commonly used low-level affect features in the literature are 68 2D facial landmark coordinates [7] and facial action unit intensity levels [2]. With the advent and success of deep representation learning, various deep learning approaches have been explored for facial affect recognition tasks [9, 18, 28, 14, 30, 11]. Compared to deep feature embeddings, LLDs are more easily interpretable and grounded in known facial expressive cues. Also, as the dimensionality of LLDs is relatively low and they are not learned using emotion labels, they are more data-efficient than deep visual features.
@@
We chose to use LLDs as facial affect descriptors in xTrace, particularly 2D facial landmark features, their confidence scores, and AU intensities over deep visual features for the following reasons:
@@
Explainability and Interpretability. Unlike deep visual features, LLDs are easy to explain and interpret as they are standardised affect descriptors grounded in visible facial appearance changes. For instance, AU intensity features capture atomic movements of facial muscles and their temporal dynamics. When used as input features, AUs render transparency which deep visual features lack due to their black-box and hierarchically compositional nature. For example, humans easily interpret the meaning of specific facial actions such as a brow raise (AU1, AU2) or a smile (AU12), whereas deep visual features often lack such direct human interpretability. Given this advantage with LLDs, we can trace the affect recognition model’s predictions back to its corresponding input features, and attribute them to visible appearance changes in different face regions. This traceability is particularly helpful in characterising the model’s failure cases and improving the model’s accuracy and robustness.
@@
Efficient use of labelled affect data. When training an end-to-end model for affect recognition from face videos, the labelled data are used for supervision for two subtasks: learning per-frame visual features from face image pixels and learning temporal dynamics that are relevant to affect prediction. The former task involves distilling facial expression visual features that are robust to nuisance factors such as changes in head pose, lighting, occlusions, etc. To achieve such a high degree of robustness, it requires large amounts of affect-labelled data or sophisticated pre-training techniques (e.g., SwaV [1]) or joint-learning methods (e.g., EmoNet [30]). To achieve high affect recognition accuracy, it also requires large amounts of temporal variations in affect labels in training sets, to learn temporal dynamics of the facial features. Achieving both these objectives requires enormously large amounts of affect-labelled data, obtaining which is prohibitively expensive due to the requirement of trained human annotators for this task. To efficiently use the labelled data, in this work we chose LLDs whose models are trained to be robust to changes in head-pose angles, lighting, and occlusions. Thus, in the proposed system all the affect-labelled data gets mainly used for learning the temporal dynamics of affect labels.
@@
Agnostic to low-level image characteristics. This is a key requirement to make an affect recognition system invariant to changes in the image acquisition process. For example, an end-to-end model trained on RGB images is likely to fail when applied to Near-Infrared (NIR) images of arbitrary wavelengths. This could be due to the dependence of deep visual features on low-level pixel statistics. Due to this limitation, an additional model fine-tuning effort may be required when there is a change in the image acquisition steps. In xTrace, by making the models that extract LLDs robust to such low-level pixel patterns, the affect recognition model is made agnostic to any low-level image characteristics.
II-B Affect Regression Outputs
@@
Time- and value-continuous predictions of affect. The affect regression model in xTrace is a sequence-to-sequence mapping model that takes a sequence of LLDs per frame as input and outputs a sequence of valence and arousal estimates. By mapping each frame to its corresponding continuous-valued vector in the 2D emotion space, this fine-grained approach allows us to make full use of the emotion representation capacity of the dimensional affect model. Also, by outputting per-frame affect vectors, this model captures transient and sudden changes in affective states over time. This level of spatio-temporal granularity broadens the applicability of xTrace for a broad spectrum of behavioural analysis tasks.
@@
Sampling-free predictive uncertainty estimates of affect. In addition to the point-wise affect predictions, xTrace also outputs three types of predictive uncertainty estimates of affect: Epistemic, Aleatoric, and cumulative [5]. Most importantly, unlike other existing methods [8, 26], it does so in a sampling-free manner, i.e., without requiring multiple forward passes through the model at inference time. This ability is important in obtaining uncertainty estimates with limited computational resources at inference time. Predictive uncertainties are essential for downstream applications to decide when they can or cannot trust the affect predictions. The three types of affect uncertainty output from xTrace have specific use cases. Epistemic uncertainty or model uncertainty allows the system to flag and collect out-of-distribution inputs and guide training data collection in later model update iterations. This feature is critical to continuously improve the coverage, accuracy, and robustness of the system over time. The aleatoric or data uncertainty captures the stochastic noise inherent in the input video. For example, when there is strong illumination on the face making it difficult to read the expression, the aleatoric uncertainty is expected to be high. When implemented in real-world applications, having access to this component of uncertainty can guide any potential automated interventions to improve the quality of face video. The cumulative uncertainty captures both epistemic and aleatoric components, and it can be made available to a downstream task to decide how reliable the affect predictions are for a given input.
II-C Building Blocks of xTrace
@@
Uncertainty-Aware Landmark Detection and Tracking. This module takes a video frame as input and outputs (a). a binary label for face shape validity (b). a 2D array of 68 landmark coordinates in iBUG format [22] and (c). a 1D array of landmark-wise uncertainty estimates. This module uses an edge-friendly CNN architecture with three output heads to jointly predict these three outputs. This CNN model is trained to be robust to in-the-wild facial appearances, on an internally curated face image dataset containing $\sim$650k RGB and NIR images labelled with 68 landmarks and their visibility scores. Note that the face-shape validity output is learned in an unsupervised fashion, whereas the landmark-wise uncertainty values are learned using explicit supervision. If the predicted face shape validity label is false for the current frame in the buffer, its corresponding affect features (both landmarks and AUs) are set to zeros and the next video frame in the buffer gets processed. If not, the predicted landmarks are used to preprocess the input to the AU intensity estimation module. It first extracts the face crop from the video frame using the bounding box computed from the previous frame’s landmark coordinates. In case there is no previous frame or its predicted face shape validity is False, the bounding box is obtained from a face detection model heavily optimised for low inference latency and high detection accuracy on the edge devices.
@@
AU Intensity Estimation model takes an image of similarity-transformed face crop as input and outputs a 15-dimensional AU intensity vector. The predicted AU intensity outputs are continuous values in the range [0,5]; 0 indicates no activation and 5 indicates maximum activation. Similar to the landmark detection and tracking model, this module also uses an edge-friendly CNN architecture for compute-efficient model inference. By training on an internally curated dataset that contains $\sim$2.4 million in-the-wild facial appearances, the AU model is made robust to the presence of partial occlusions and non-uniform illumination.
@@
III Video Datasets for Facial Affect Recognition